OpenStudy (anonymous):

Solve for n: 1/5n-12=-3

OpenStudy (anonymous):

isolate the N by itself... using algebra rules - how do we start?

OpenStudy (anonymous):

1/5n=-3+12 1/5n=9

OpenStudy (anonymous):

so what does n=

OpenStudy (anonymous):

you need to isolate to find that...so look at the problem. you need to get N on the side by itself

OpenStudy (anonymous):

Sorry im not so good at this stuff

OpenStudy (kc_kennylau):

Multiply both sides by 5